# Many-Shot In-Context Learning in Multimodal Foundation Models

### v1

@@ Abstract
Large language models are well-known to be effective at few-shot in-context learning (ICL). Recent advancements in multimodal foundation models have enabled unprecedentedly long context windows, presenting an opportunity to explore their capability to perform ICL with many more demonstrating examples. In this work, we evaluate the performance of multimodal foundation models scaling from few-shot to many-shot ICL. We benchmark GPT-4o and Gemini 1.5 Pro across 10 datasets spanning multiple domains (natural imagery, medical imagery, remote sensing, and molecular imagery) and tasks (multi-class, multi-label, and fine-grained classification). We observe that many-shot ICL, including up to almost 2,000 multimodal demonstrating examples, leads to substantial improvements compared to few-shot (<100 examples) ICL across all of the datasets. Further, Gemini 1.5 Pro performance continues to improve log-linearly up to the maximum number of tested examples on many datasets. Given the high inference costs associated with the long prompts required for many-shot ICL, we also explore the impact of batching multiple queries in a single API call. We show that batching up to 50 queries can lead to performance improvements under zero-shot and many–shot ICL, with substantial gains in the zero-shot setting on multiple datasets, while drastically reducing per-query cost and latency. Finally, we measure ICL data efficiency of the models, or the rate at which the models learn from more demonstrating examples. We find that while GPT-4o and Gemini 1.5 Pro achieve similar zero-shot performance across the datasets, Gemini 1.5 Pro exhibits higher ICL data efficiency than GPT-4o on most datasets. Our results suggest that many-shot ICL could enable users to efficiently adapt multimodal foundation models to new applications and domains. Our codebase is publicly available at https://github.com/stanfordmlgroup/ManyICL.
@@ 1 Introduction
Large language models (LLMs) have been shown to substantially benefit from the inclusion of a few demonstrating examples (shots) in the LLM context before the test query [1, 2, 3]. This phenomenon, commonly referred to as in-context learning (ICL), enables LLMs to learn from few shots without any updates to model parameters, and therefore improves specialization to new tasks without any further model training. More recently, large multimodal models (LMMs) have also demonstrated the capability of learning from in-context examples [4, 5, 6]. Han et al. [5] and Zhang et al. [6] both show that few-shot multimodal ICL specifically helps to improve LMM performance on out-domain or out-of-distribution tasks.
@@
To investigate the capability of state-of-the-art multimodal foundation models to perform many-shot ICL, we conduct a large suite of experiments benchmarking model performance on 10 datasets spanning several domains and image classification tasks after scaling up the number of demonstrating examples by multiple orders of magnitude. Specifically, our contributions are as follows:
@@
    We show that providing multimodal foundation models with many demonstrating examples leads to substantial performance improvements compared to providing only a few demonstrating examples. We observe that the performance of Gemini 1.5 Pro generally improves log-linearly as the number of demonstrating examples increases, whereas GPT-4o exhibits less stable improvements as the number of in-context examples increases.
@@
    We measure the data efficiency of the models under ICL as the number of demonstrating examples is increased, and find that Gemini 1.5 Pro exhibits higher ICL data efficiency than GPT-4o on most datasets.
@@
Our codebase is publicly available at https://github.com/stanfordmlgroup/ManyICL.
@@ 2 Related Work
Scaling ICL.  The seminal work of Brown et al. [1] discovered performance improvements for LLMs from increasing the number of in-context examples, but the tested number of demonstrating examples was low (10 to 100), likely due to the restrictive context size (2048 tokens for GPT3). Increasing the number of in-context examples has only been explored recently by a few works [7, 8, 9]. Both Li et al. [7] and Agarwal et al. [8] explore scaling in-context learning to more than 1,000 demonstrating examples and find performance improvements across multiple tasks. However, their experiments are limited to text-only benchmarks and do not compare performance across different models.
Multimodal ICL.  Due to the recent emergence of LMMs, research on multimodal ICL is still nascent. One prior work developed a new model to leverage complex prompts composed of multimodal inputs in order to allow models to compare images [10], while other recent works explored the generalizability of GPT-4V and Gemini to multimodal out-domain and out-of-distribution tasks, and found that ICL leads to performance benefits for both models across many tasks [6, 5]. However, none of these works have leveraged the new largely expanded context windows to investigate the effects of increasing the number of demonstrating examples.
Batch Querying.  Multiple prior works have explored batching queries (also commonly referred to as batch prompting) for more efficient and cheaper inference. Batch prompting was first introduced in Cheng et al. [11], leading to comparable or better performance than single prompting, while achieving substantially reduced inference token cost and latency. Lin et al. [12] observe performance degradation with batched prompts in longer contexts, and propose a variety of techniques to mitigate the performance loss. More recently, additional variations of batch prompting have been proposed, including grouping similar questions together [13], batching prompts of different tasks [14], and concatenating multiple images into a single image collage [15]. We again note that batch prompting with high numbers of demonstrating examples and high numbers of queries has only become feasible due to larger context windows of recent models.
@@ 3 Methods
Figure 2: Gemini 1.5 Pro and GPT-4o performance from zero-shot to many-shot ICL. X-axis is in log scale. For Gemini 1.5 Pro, we observe log-linear improvement on 9 out of the 10 datasets and for GPT-4o we observe improvement from more demonstrating examples on most datasets, albeit substantially less stable than Gemini 1.5 Pro.
We conduct several experiments to test the effect of increasing the number of demonstrating examples on the performance of two state-of-the-art multimodal foundation models: GPT-4o and Gemini 1.5 Pro (Section 3.1). We benchmark their performance using standard performance metrics as well as an ICL data efficiency metric (Section 3.3) on 10 datasets spanning several vision domains and image classification tasks (Section 3.2). We conduct ablation studies to test the impact of batching queries on model performance and explain the substantial improvement in zero-shot settings (Section 4.2). We refer to the many-shot in-context learning framework as many-shot ICL. Figure 1 provides an illustrative summary of many-shot ICL and batched many-shot ICL compared to zero-shot and few-shot ICL.
@@ 3.1 Models
We use three state-of-the-art multimodal foundation models with public API access, namely GPT-4o, GPT4(V)-Turbo [4], and Gemini 1.5 Pro [16]. Because GPT-4o performs substantially better than GPT4(V)-Turbo, we focus on the results of GPT-4o and Gemini 1.5 Pro in the main text, and include GPT4(V)-Turbo results in the Appendix. We do not utilize Claude3-Opus in our experiments, as it only accepts up to 20 images in one request at the time of writing. The specific endpoint for for GPT-4o is “gpt-4o-2024-05-13”, for GPT-4(V)-Turbo is “gpt-4-turbo-2024-04-09”, and for Gemini 1.5 Pro is “gemini-1.5-pro-preview-0409”. We use the API service provided by OpenAI for GPT-4o and GPT-4(V)-Turbo, and the API service provided by Google Cloud on Vertex AI for Gemini 1.5 Pro. We set the temperature to zero for all models and a random seed for GPT-4(V)-Turbo and GPT-4o to obtain more deterministic responses. To prevent models from abstaining (which happens rarely), we rerun the query until an answer is provided.
@@ 3.2 Datasets
We benchmark the model performance on 10 datasets spanning multiple domains (natural imagery, medical imagery, remote sensing, and molecular imagery) and tasks (multi-class, multi-label, and fine-grained classification). We choose to focus on image classification tasks as other tasks such as region captioning would require substantially more tokens thereby limiting the total number of demonstrating examples, and most LMMs are not yet capable of accurately producing localizations required for other tasks like bounding boxes and segmentation masks [17, 18]. Table 1 provides a summary of the datasets used in this study.
For all datasets, we construct a set of demonstration (demo) examples from the original training and validation splits used for in-context learning and a test set from the original test split (if one exists) to evaluate the performance of the models. We randomly sample the demo and test sets from the original dataset without replacement. For the multi-class and fine-grained classification datasets, we perform a class-stratified sampling, ensuring an equal number of examples per class in both the demo and test sets. For the multi-label classification dataset (CheXpert), we sample an equal number of positive and negative samples per class in both the demo and test sets. We note that, since the task is multi-label, this sampling procedure does not result in an exactly equal number of examples per class. The per-dataset sizes of the full demo and test sets are shown in Table 1, and we increase the number of demonstration examples up to the numbers shown in the table while ensuring class balance for the scaling experiments.
Table 1: Summary of benchmark datasets. We use 10 datasets spanning multiple domains (natural imagery, medical imagery, remote sensing, molecular imagery) and tasks (multi-class, multi-label, and fine-grained classification).
Dataset Task and image type # Classes Demo / test set size Example image
HAM10000[19] Skin disease classification on clinical photos 7 805 / 210 [Uncaptioned image]
FIVES [20] Eye disease classification on fundus images 4 400 / 120 [Uncaptioned image]
CheXpert [21] Multi-label lung disease detection on chest X-rays 5 200 / 150 [Uncaptioned image]
Camelyon17 [22] Tumor detection on pathology images 2 2000 / 100 [Uncaptioned image]
TerraIncognita [23] Animal species recognition on camera images 9 1035 / 270 [Uncaptioned image]
UCMerced[24] Land use classification on satellite images 21 1470 / 420 [Uncaptioned image]
EuroSAT [25] Land use / land cover classification on satellite images 10 1000 / 300 [Uncaptioned image]
Oxford Pets [26] Pet classification on camera images 35 1750 / 700 [Uncaptioned image]
DTD [27] Texture classification on synthetic images 47 2350 / 940 [Uncaptioned image]
DrugOOD Assay [28] Drug binding prediction on molecular images 2 1600 / 200 [Uncaptioned image]
@@ 3.3 Evaluation Metrics
We use standard metrics to evaluate model performance on each dataset. Specifically, we measure performance using accuracy for all multi-class classification datasets as they are sampled to have a balanced class distribution. For multi-label classification on CheXpert, we use the macro-averaged F1 metric. In the rare case of parsing errors, we consider the response as incorrect. To estimate the variability around the evaluation metrics, we compute standard deviation using bootstrapping with 1,000 bootstrap replicates.
@@ 4.1 Increasing number of demonstrating examples
Main Results. Gemini 1.5 Pro exhibits consistent and substantial improvements as the number of demonstrating examples increases across all datasets except for DrugOOD Assay (Figure 2). Gemini 1.5 Pro shows particularly large improvements from many-shot ICL on HAM10000 (+23% accuracy compared to zero-shot, +16% compared to 7 examples), FIVES (+29% compared to zero-shot, +27% compared to 20 examples), and EuroSAT (+38% compared to zero-shot, +31% compared to 10 examples). Notably, for 5 out of the 10 datasets (FIVES, UCMerced, EuroSAT, Oxford Pets, and DTD), Gemini 1.5 Pro performance continues to improve up to the highest number of demonstrating examples considered (~1,000 examples). On the other 5 datasets, the optimal performance occurs prior to the highest number of demo examples, with the maximum number of demo examples leading to similar or slightly worse performance than the optimal demo set size. On the other hand, Gemini 1.5 Pro performance on DrugOOD Assay does not substantially benefit from many-shot ICL, with high variance in performance across demo sizes and the peak performance at 40 demo examples.
@@
Table 2: Many-shot ICL performance and efficiency comparison. We report the performance under a zero-shot regime and performance at the optimal demo set size as well as the many-shot ICL data efficiency of GPT-4o and Gemini 1.5 Pro. We measure performance using accuracy on all datasets except CheXpert, for which we use macro-average F1. We bold the highest ICL data efficiency between the two models on each dataset.
@@
Figure 3: Gemini 1.5 Pro performance under many-shot and zero-shot ICL when varying the amount of queries included in every request. We show performance per batch size with the optimal number of demo examples (many-shot) and no demo examples (zero-shot). The $x$-axis is in log scale. Under the many-shot regime, batching queries leads to no substantial drop in performance compared to individual queries when we choose a suitable batch size. For zero-shot, including only one query is suboptimal for many datasets.

### v2

@@ Abstract
Large language models are well-known to be effective at few-shot in-context learning (ICL). Recent advancements in multimodal foundation models have enabled unprecedentedly long context windows, presenting an opportunity to explore their capability to perform ICL with many more demonstrating examples. In this work, we evaluate the performance of multimodal foundation models scaling from few-shot to many-shot ICL. We benchmark GPT-4o and Gemini 1.5 Pro across 14 datasets spanning multiple domains (natural imagery, medical imagery, remote sensing, and molecular imagery) and tasks (image classification, visual question answering, and object localization). We observe that many-shot ICL, including up to almost 2,000 multimodal demonstrating examples, leads to substantial improvements compared to few-shot ($<$100 examples) ICL across all of the datasets. Further, Gemini 1.5 Pro performance continues to improve log-linearly up to the maximum number of tested examples on many datasets. We also find open-weights multimodal foundation models like Llama 3.2-Vision and InternLM-XComposer2.5 do not benefit from the demonstrating examples, highlighting an important gap between open and closed multimodal foundation models. Given the high inference costs associated with the long prompts required for many-shot ICL, we also explore the impact of batching multiple queries in a single API call. We show that batching up to 50 queries can lead to performance improvements under zero-shot and many–shot ICL, with substantial gains in the zero-shot setting on multiple datasets, while drastically reducing per-query cost and latency. Finally, we measure ICL data efficiency of the models, or the rate at which the models learn from more demonstrating examples. We find that while GPT-4o and Gemini 1.5 Pro achieve similar zero-shot performance across the datasets, Gemini 1.5 Pro exhibits higher ICL data efficiency than GPT-4o on most datasets. Our results suggest that many-shot ICL could enable users to efficiently adapt multimodal foundation models to new applications and domains.
@@ 1 Introduction
Large language models (LLMs) have been shown to substantially benefit from the inclusion of a few demonstrating examples (shots) in the LLM context before the test query (Brown et al., 2020; Parnami & Lee, 2022; Wang et al., 2020). This phenomenon, commonly referred to as in-context learning (ICL), enables LLMs to learn from few shots without any updates to model parameters, and therefore improves specialization to new tasks without any further model training. More recently, large multimodal models (LMMs) have also demonstrated the capability of learning from in-context examples Achiam et al. (2023); Han et al. (2023); Zhang et al. (2024b). Han et al. (2023) and Zhang et al. (2024b) both show that few-shot multimodal ICL specifically helps to improve LMM performance on out-domain or out-of-distribution tasks.
@@
To investigate the capability of state-of-the-art multimodal foundation models to perform many-shot ICL, we conduct a large suite of experiments benchmarking model performance on 14 datasets spanning several domains and multimodal tasks after scaling up the number of demonstrating examples by multiple orders of magnitude. Specifically, our contributions are as follows:
@@
    We show that providing close-weights multimodal foundation models with many demonstrating examples leads to substantial performance improvements compared to providing only a few demonstrating examples. We observe that the performance of Gemini 1.5 Pro generally improves log-linearly as the number of demonstrating examples increases, whereas GPT-4o exhibits less stable improvements as the number of in-context examples increases.
@@
    We find open-weights multimodal foundation models like Llama 3.2-Vision and InternLM-XComposer2.5 do not benefit from the demonstrating examples, highlighting a significant gap and an important direction for the open-weights community.
@@
    We measure the data efficiency of the models under ICL as the number of demonstrating examples increases, and find that Gemini 1.5 Pro exhibits higher ICL data efficiency than GPT-4o on most datasets.
  4. 4.
@@
  6. 6.
    We release a new tool (anonymized link) to allow users to easily test the many-shot ICL capabilities of multimodal foundation models, facilitating future work on studying them.
@@ 2 Related Work
Scaling ICL.  The seminal work of Brown et al. (2020) discovered performance improvements for LLMs from increasing the number of in-context examples, but the tested number of demonstrating examples was low (10 to 100), likely due to the restrictive context size (2048 tokens for GPT3). Increasing the number of in-context examples has only been explored recently by a few works Li et al. (2023); Agarwal et al. (2024); Bertsch et al. (2024). Both Li et al. (2023) and Agarwal et al. (2024) explore scaling in-context learning to more than 1,000 demonstrating examples and find performance improvements across multiple tasks. However, their experiments are limited to text-only benchmarks and do not compare performance across different models.
Multimodal ICL.  Due to the recent emergence of LMMs, research on multimodal ICL is still nascent. One prior work developed a new model to leverage complex prompts composed of multimodal inputs in order to allow models to compare images Zhao et al. (2023), while other recent works explored the generalizability of GPT-4V and Gemini to multimodal out-domain and out-of-distribution tasks, and found that ICL leads to performance benefits for both models across many tasks Zhang et al. (2024b); Han et al. (2023). However, none of these works have leveraged the new largely expanded context windows to investigate the effects of increasing the number of demonstrating examples.
Batch Querying.  Multiple prior works have explored batching queries (also commonly referred to as batch prompting) for more efficient and cheaper inference. Batch prompting was first introduced in Cheng et al. (2023), leading to comparable or better performance than single prompting, while achieving substantially reduced inference token cost and latency. Lin et al. (2023) observe performance degradation with batched prompts in longer contexts, and propose a variety of techniques to mitigate the performance loss. More recently, additional variations of batch prompting have been proposed, including grouping similar questions together Liu et al. (2024), batching prompts of different tasks Son et al. (2024), and concatenating multiple images into a single image collage Xu et al. (2024). We again note that batch prompting with high numbers of demonstrating examples and high numbers of queries has only become feasible due to larger context windows of recent models.
@@ 3 Methods
Figure 2: Gemini 1.5 Pro and GPT-4o performance from zero-shot to many-shot ICL. The x-axis is in log scale. For Gemini 1.5 Pro, we observe log-linear improvement on 9 out of the 10 datasets and. For GPT-4o, we observe improvement from more demonstrating examples on most datasets, while the improvement is substantially less stable than Gemini 1.5 Pro. Error bars are estimated standard deviations using bootstrapping with 1,000 bootstrap replicates.
Table 1: Summary of datasets. We use 14 datasets spanning multiple domains (natural imagery, medical imagery, remote sensing, and molecular imagery) and tasks (multi-class, multi-label, and fine-grained classification, visual question answering, object localization).
Dataset Task and image type Demo / test set size Example image
HAM10000
(Tschandl et al., 2018)
7-category skin disease classification on clinical photos 805 / 210 [Uncaptioned image]
FIVES
(Jin et al., 2022)
4-category eye disease classification on fundus images 400 / 120 [Uncaptioned image]
CheXpert
(Irvin et al., 2019)
Multi-label 5-category lung disease detection on chest X-rays 200 / 150 [Uncaptioned image]
Camelyon17
(Bandi et al., 2018)
Binary tumor detection on pathology images 2000 / 100 [Uncaptioned image]
TerraIncognita
(Beery et al., 2018)
9-category animal species recognition on camera images 1035 / 270 [Uncaptioned image]
UCMerced
(Yang & Newsam, 2010)
21-category land use classification on satellite images 1470 / 420 [Uncaptioned image]
EuroSAT
(Helber et al., 2019)
10-category land use / land cover classification on satellite images 1000 / 300 [Uncaptioned image]
Oxford Pets
(Parkhi et al., 2012)
35-category pet classification on camera images 1750 / 700 [Uncaptioned image]
DTD
(Cimpoi et al., 2014)
47-category texture classification on synthetic images 2350 / 940 [Uncaptioned image]
DrugOOD Assay
(Ji et al., 2022)
Binary drug binding prediction on molecular images 1600 / 200 [Uncaptioned image]
RSVQA
(Lobry et al., 2020)
Visual question answering on satellite images 200/200 [Uncaptioned image]
VQA-RAD
(Lau et al., 2018)
Visual question answering on radiology images 200/200 [Uncaptioned image]
DIOR
(Li et al., 2020)
Object localization on satellite images 200/100 [Uncaptioned image]
DeepLesion
(Yan et al., 2018)
Lesion localization on CT images 200/100 [Uncaptioned image]
We conduct several experiments to test the effect of increasing the number of demonstrating examples on the performance of state-of-the-art multimodal foundation models: close-weights models like GPT-4o and Gemini 1.5 Pro, and open-weights models like Llama3.2-Vision (Section 3.1). We benchmark their performance using standard performance metrics as well as an ICL data efficiency metric (Section 3.3) on 14 datasets spanning several vision domains and multimodal tasks (Section 3.2). We conduct ablation studies to test the impact of batching queries on model performance and explain the substantial improvement in zero-shot settings (Section 4.2). We refer to the many-shot in-context learning framework as many-shot ICL. Figure 1 provides an illustrative summary of many-shot ICL and batched many-shot ICL compared to zero-shot and few-shot ICL.
@@ 3.1 Models
We use three state-of-the-art multimodal foundation models with public API access, namely GPT-4o, GPT4(V)-Turbo (Achiam et al., 2023), and Gemini 1.5 Pro (Reid et al., 2024). In addition, two open-weights multimodal foundation models (Llama3.2-11B-Vision (lla, ) and InternLM-XComposer-2.5 (Zhang et al., 2024a)) are also tested. Because GPT-4o performs substantially better than GPT4(V)-Turbo, we focus on the results of GPT-4o and Gemini 1.5 Pro in the main text, and include GPT4(V)-Turbo results in the Appendix. We do not utilize Claude3-Opus in our experiments, as it only accepts up to 20 images in one request at the time of writing. The specific endpoint for for GPT-4o is “gpt-4o-2024-05-13”, for GPT-4(V)-Turbo is “gpt-4-turbo-2024-04-09”, and for Gemini 1.5 Pro is “gemini-1.5-pro-preview-0409”. We use the API service provided by OpenAI for GPT-4o and GPT-4(V)-Turbo, and the API service provided by Google Cloud on Vertex AI for Gemini 1.5 Pro. We set the temperature to zero for all models and a random seed for GPT-4(V)-Turbo and GPT-4o to obtain more deterministic responses. To prevent models from abstaining (which happens rarely), we rerun the query until an answer is provided. We call the APIs on a virtual machine instance of type “c2-standard-8” and run inference using open-weights models on a “a3-highgpu-8g” machine (with 8 H100 GPUs) hosted on Google Cloud Platform.
@@ 3.2 Datasets
We benchmark the model performance on 14 datasets spanning multiple domains (natural imagery, medical imagery, remote sensing, and molecular imagery) and tasks (image classification, visual question answering, and object localization). We acknowledge most LMMs are not yet capable of accurately producing localizations (Wu et al., 2024; Zang et al., 2023). Table 1 provides a summary of the datasets used in this study.
For all datasets, we construct a set of demonstrating (demo) examples from the original training and validation splits used for in-context learning and a test set from the original test split (if one exists) to evaluate the performance of the models. We randomly sample the demo and test sets from the original dataset without replacement. For the multi-class and fine-grained classification datasets, we perform a class-stratified sampling, ensuring an equal number of examples per class in both the demo and test sets. For the multi-label classification dataset (CheXpert), we sample an equal number of positive and negative samples per class in both the demo and test sets. We note that, since the task is multi-label, this sampling procedure does not result in an exactly equal number of examples per class. For the scaling experiments, we increase the number of demonstrating examples while ensuring class balance. Besides the 10 classification datasets shown in Table 1, we also include two visual question answering datasets (RSVQA (Lobry et al., 2020) and VQA-RAD (Lau et al., 2018)) and two object localization datasets (DIOR for bridge localization (Li et al., 2020) and DeepLesion for lung lesion localization (Yan et al., 2018)).
@@ 3.3 Evaluation Metrics
We use standard metrics to evaluate model performance on each dataset. Specifically, we measure performance using accuracy for all multi-class classification and visual question answering datasets as they are sampled to have a balanced class distribution. For multi-label classification on CheXpert, we use the macro-averaged F1 metric. In the rare case of parsing errors, we consider the response as incorrect. For object localization datasets, we use mean Intersection over Union(IoU). To estimate the variability around the evaluation metrics, we compute standard deviation using bootstrapping with 1,000 bootstrap replicates, and it captures the variability in data sampling.
@@
Figure 3: Llama3.2 and InternLM-XComposer2.5 performance compared with Gemini 1.5 Pro. While Gemini 1.5 Pro shows substantial improvement from zero-shot to many-shot ICL on all three categories of tasks, both open-weights models do not benefit from demonstrating examples.
@@ 4.1 Increasing number of demonstrating examples
Main Results. Gemini 1.5 Pro exhibits consistent and substantial improvements as the number of demonstrating examples increases across all datasets except for DrugOOD Assay (Figure 2 and  3). Gemini 1.5 Pro shows particularly large improvements from many-shot ICL on HAM10000 (+23% accuracy compared to zero-shot, +16% compared to 7 examples), FIVES (+29% compared to zero-shot, +27% compared to 20 examples), and EuroSAT (+38% compared to zero-shot, +31% compared to 10 examples). Notably, for 8 out of the 14 datasets (FIVES, UCMerced, EuroSAT, Oxford Pets, DTD, VQA-RAD, DIOR and DeepLesion), Gemini 1.5 Pro performance continues to improve up to the highest number of demonstrating examples considered (~1,000 examples). On the other 6 datasets, the optimal performance occurs prior to the highest number of demo examples, with the maximum number of demo examples leading to similar or slightly worse performance than the optimal demo set size. On the other hand, Gemini 1.5 Pro performance on DrugOOD Assay does not substantially benefit from many-shot ICL, with high variance in performance across demo sizes and the peak performance at 40 demo examples.
@@
Open-weights Model Results. We also find open-weights multimodal foundation models like Llama 3.2-Vision and InternLM-XComposer2.5 do not benefit from the demonstration examples (Figure 3), highlighting a significant gap between open and closed multimodal foundation models.
@@
Table 2: Many-shot ICL performance and efficiency comparison. We report the performance under a zero-shot regime and performance at the optimal demo set size as well as the many-shot ICL data efficiency of GPT-4o and Gemini 1.5 Pro. We measure performance using accuracy on all datasets except CheXpert, for which we use macro-averaged F1. The highest ICL data efficiency between the two models on each dataset is marked in bold.